OpenStudy (anonymous):

Solve the system of equations: what is the y-value in the solution? 3x - 5y = -16 and 2x + 5y = 31 A. 5 B. 3 C. 2 D. 4

hero (hero):

3x - 5y = -16 2x + 5y = 31 5x = 15 x = 15/5 x = 3 y = 5 The solution is A, 5.